Q. F( x + 2y, x – 2y) = xy , then F(X,Y) EQUALS- x 2 - y 2 /8 x 2 – y 2 /4 x 2 + y 2 /4 x 2 – y 2 /2

Q.  F( x + 2y, x – 2y) = xy , then F(X,Y) EQUALS-
  1. x- y2/8
  2. x2 – y2/4
  3. x2 + y2/4
  4. x2 – y2/2

Grade:Select Grade

1 Answers

Y RAJYALAKSHMI
9 years ago
Ans: 
  1. x- y2/8
 
if, F(xy) = x2 – y2/8, then F( x + 2y, x – 2y) = (x+2y)2 – (x-2y)2/8 
= 4*x*2y/8 = xy
 
Hence ans is 
  1. x- y2/8